# Hulls of projective Reed-Muller codes over the projective plane

### v2

@@ 3.2. Hermitian hull
We have $f(0,0,1)=0$, which means that $\mu_{0,0,d^{\perp}}=0$. Following the proof of Lemma 3.5, we consider $\operatorname{in}(f)=x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}\bmod I(P^{2})$, with $a_{1}>0$, $10\leq a_{2}\leq q^{2}-$ 1 and $a_{1}+a_{2}=d$. Because of (9) we must have some monomial in $(A^{d^{\perp}})^{q}$ such that its expression in the basis from Lemma 2.10 contains $x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}$ in its support. Let $\gamma_{2}=\overline{qa_{2}}$, which implies that $q\gamma_{2}\equiv a_{2}\bmod q^{2}-1$. The only monomials in $(A^{d^{\perp}})^{q}$ that contain $x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}$ in their expression in terms of the basis from Lemma 2.10 are $(x_{1}^{d^{\perp}-\gamma_{2}}x_{2}^{\gamma_{2}})^{q}$ if $q(d^{\perp}-\gamma_{2})\equiv a_{1}\bmod q^{2}-1$, and $(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}$ with $q\gamma_{1}\equiv a_{1}\bmod q^{2}-1$, $11\leq\gamma_{1}\leq q^{2}-$ 1 (i.e., $\gamma_{1}=\overline{qa_{1}}$), if $\gamma_{0}=d^{\perp}-\gamma_{1}-\gamma_{2}>0$.
In the first case, $q(d^{\perp}-\gamma_{2})\equiv a_{1}\bmod q^{2}-1$ implies that $qd^{\perp}\equiv d\bmod q^{2}-1$, which happens if and only if $d\equiv 0\bmod q-1$. Taking into account that $d\leq 2(q-1)$, we have $d^{\perp}>q^{2}-1\geq\gamma_{2}$, and we have $x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}\equiv(x_{1}^{d^{\perp}-\gamma_{2}}x_{2}^{\gamma_{2}})^{q}\bmod I(P^{2})$. Moreover, in this situation we can do this for all the monomials from $A_{2}^{d}$ in the support of $f$.
If $d\not\equiv 0\bmod q-1$, the only monomial in $(A^{d^{\perp}})^{q}$ with $x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}$ in its expression in terms of the elements of the basis from Lemma 2.10 is $(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}$ with $\gamma_{0}=d^{\perp}-\gamma_{1}-\gamma_{2}$, if $d^{\perp}-\gamma_{1}-\gamma_{2}>0$. This is because, by Lemma 2.11, we have that
$(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}\equiv x_{1}^{\overline{a_{1}}}x_{2}^{a_{2}}+(x_{0}-1)(x_{2}^{a_{2}}+x_{1}-1)\bmod I(P^{2}),$
@@
Arguing as in Lemma 3.5 and using Lemma 3.16, we obtain that there must be a sum of monomials in $((A^{d^{\perp}})^{q})^{(l)}$ equal to $\lambda_{0,d}\left((x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}\right)$. This implies that we have the evaluation of $(x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}$ in $\operatorname{PRM}_{d}^{\perp_{h}}(2)$. This is a contradiction because we have the evaluation of $(x_{0}^{d^{\perp}-\beta_{2}}x_{2}^{\beta_{2}})^{q}$ in $\operatorname{PRM}_{d}^{\perp_{h}}(2)$, and $(x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}-(x_{0}^{d^{\perp}-\beta_{2}}x_{2}^{\beta_{2}})^{q}\equiv(x_{0}-1)(x_{1}-1)\bmod I(P^{2})$, whose evaluation has Hamming weight 1. ∎
@@
To state the next result, we use the following sets of polynomials. We recall that $U=\{x_{0}^{d-a_{1}-a_{2}}x_{1}^{a_{1}}x_{2}^{a_{2}}\mid x_{1}^{a_{1}}x_{2}^{a_{2}}\in U_{d-1,d}\}$, where we consider $U_{d-1,d}$ as in Proposition 3.14. We define $V:=\{x_{1}^{d-a_{2}}x_{2}^{a_{2}}\mid a_{2}\in T\}$, where $T$ is defined as in Lemma 3.19. Finally, for the case $12(q-1)<d\leq q^{2}-$ 1 we define $W:=\{x_{1}^{d-a_{2}}x_{2}^{a_{2}}+x_{0}^{d-\overline{qd^{\perp}-a_{2}}-a_{2}}x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}\mid q^{2}-1\geq d^{\perp}-\overline{qa_{2}}>\overline{q(d-a_{2})}\text{ and }d-a_{2}>\overline{qd^{\perp}-a_{2}}\}$. We are interested in $W$ because, if $d^{\perp}-\overline{qa_{2}}>\overline{q(d-a_{2})}$ and $d-a_{2}>\overline{qd^{\perp}-a_{2}}$, then
@@
This equivalence can be checked by considering the evaluation of both polynomials at $P^{2}$. Therefore, we have $W\subset S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. The previous equivalence explains the definition of $W$, except the condition $q^{2}-1\geq d^{\perp}-\overline{qa_{2}}$. This conditions ensures that the set $U\cup V\cup W$ is linearly independent in $S/I(P^{2})$. Indeed, both of the monomials in the left hand side of (16) are monomials from the basis of Lemma 2.8 and are not in $U\cup V$. The monomial $x_{1}^{d-a_{2}}x_{2}^{a_{2}}$ from (16) is not in $V$ because $d^{\perp}\leq\overline{qa_{2}}-(q^{2}-1)$ (see the definition of $T$ in Lemma 3.19).
For the monomial $x_{0}^{d-\overline{qd^{\perp}-a_{2}}-a_{2}}x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}$, this monomial is not in $U$ because we have $x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}\not\in U_{d-1,d}$. To see this, we use the definition of $U_{d-1,d}$ from Proposition 3.14. First, $d-a_{2}>\overline{qd^{\perp}-a_{2}}$ means that this monomial satisfies the first condition in the definition of $U_{d-1,d}$. For the second one, we would have to check if
(17) $\overline{q(qd^{\perp}-a_{2})}+\overline{qa_{2}}=\overline{d^{\perp}-\overline{qa_{2}}}-\overline{qa_{2}}\leq d^{\perp}-1.$
If $\overline{d^{\perp}-\overline{qa_{2}}}<d^{\perp}-\overline{qa_{2}}$, we clearly have (17). However, if $\overline{d^{\perp}-\overline{qa_{2}}}=d^{\perp}-\overline{qa_{2}}$, we do not have (17), and this happens if and only if $q^{2}-1\geq d^{\perp}-\overline{qa_{2}}$, which is the condition we are using in the definition of $W$. With all of this, we can state the main result of this section.
@@
Let $11\leq d\leq q^{2}-$ 1, and $U=\{x_{0}^{d-a_{1}-a_{2}}x_{1}^{a_{1}}x_{2}^{a_{2}}\mid x_{1}^{a_{1}}x_{2}^{a_{2}}\in U_{d-1,d}\}$, where we consider $U_{d-1,d}$ as in Proposition 3.14. Let $V$ and $W$ be as in the discussion above. If $d\equiv 0\bmod q-1$, then $U\cup A_{2}^{d}\cup A_{3}^{d}$ is a basis for $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. For $d\not\equiv 0\bmod q-1$, if $d\leq 2(q-1)$, then $U\cup V=A_{1}^{d}\cup V$ is a basis for $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Lastly, if $12(q-1)<d\leq q^{2}-$ 1, then $U\cup V\cup W$ is a linearly independent set contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$.
@@
If $d\equiv 0\bmod q-1$ and $d\leq q^{2}-1$, reasoning as in the proofs of Lemmas 3.16, 3.19 and 3.20, we have $U\cup A_{2}^{d}\cup A_{3}^{d}$ contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Let $f\in S_{d}$ such that $f\in S_{d^{\perp}}^{q}/I(P^{2})$ and whose support is contained in $A_{1}^{d}\setminus U$. Then $f$ has an expression in terms of the monomials from $(A^{d^{\perp}})^{q}$ in $S/I(P^{2})$, and this expression only involves the monomials from $(A_{1}^{d^{\perp}})^{q}$. This is because $x_{2}^{qd^{\perp}}$ cannot be in the expression because it is nonzero at $[0:0:1]$ while $f(0,0,1)=0$, and if there were monomials from $(A_{2}^{d^{\perp}})^{q}$, in [$[\{0\}\times\{1\}\times{\mathbb{F}}_{q^{2}}]$ that expression would have the same evaluation as some polynomial in $x_{2}$ with degree less than or equal to $q^{2}-1$, which cannot have $q^{2}$ zeroes ($f$ is equal to 00 in those points). We finish this case by using the affine case from Proposition 3.14.
@@
Finally, in the case $d\not\equiv 0\bmod q-1$ and $12(q-1)<d<q^{2}-$ 1, we have from Lemma 3.16 that $U$ is contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. The fact that $V$ is contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ follows from (12). For $W$, we saw in the discussion above this result that it is also contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$, and we also obtained that $U\cup V\cup W$ is linearly independent in $S/I(P^{2})$. ∎
From Theorem 3.21 we can obtain the exact dimension of $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ if $d\leq 2(q-1)$ or $d\equiv 0\bmod q-1$. For $12(q-1)<d\leq q^{2}-$ 1, $d\not\equiv 0\bmod q-1$, we only have a lower bound for the dimension of $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Note that a lower bound for the dimension of the hull gives an upper bound for the parameter $c$ of the corresponding EAQECC, which is still interesting because it tells us how many maximally entangled pairs are required at most for using that EAQECC. Equivalently, if we use as many maximally entangled pairs as the bound specifies, then we can employ this EAQECC. Nevertheless, in all cases we have checked, this is indeed the true value of the dimension of the hull, which implies that $U\cup V\cup W$ is also a basis for the hull in those cases. In particular, this means that the Hermitian hull is not generated by monomials in general because of $W$ (we saw in the discussion prior to Theorem 3.21 that no monomial of the polynomials from $W$ is contained in $U\cup V$). We see this in the next example.
@@ 4.2. Hermitian EAQECCs
Let $11\leq d<q^{2}-$ 1. Then we can construct an EAQECC with parameters $[[n,\kappa,\delta;c]]_{q}$, where $n=q^{4}$, $\kappa=n-2(\dim\operatorname{RM}_{d}(q^{2},2))+c$, $\delta\geq\operatorname{wt}(\operatorname{RM}_{d}^{\perp}(q^{2},2))$, and the value of $c$ is

### v3

@@ 3.2. Hermitian hull
We have $f(0,0,1)=0$, which means that $\mu_{0,0,d^{\perp}}=0$. Following the proof of Lemma 3.5, we consider $\operatorname{in}(f)=x_{1}^{a_{1}}x_{2}^{a_{2}}$, with $a_{1}>0$, $10\leq a_{2}\leq q^{2}-$ 1 and $a_{1}+a_{2}=d$ (since $d\leq 2(q-1)$, we also have $a_{1}\leq q^{2}-1$). Because of (9) we must have some monomial in $(A^{d^{\perp}})^{q}$ such that its expression in the basis from Lemma 2.9 contains $x_{1}^{a_{1}}x_{2}^{a_{2}}$ in its support. Let $\gamma_{i}=\overline{qa_{i}}$, for $i=1,2$, which implies that $q\gamma_{i}\equiv a_{i}\bmod q^{2}-1$. The only monomials in $(A^{d^{\perp}})^{q}$ that contain $x_{1}^{a_{1}}x_{2}^{a_{2}}$ in their expression in terms of the basis from Lemma 2.9 are $(x_{1}^{d^{\perp}-\gamma_{2}}x_{2}^{\gamma_{2}})^{q}$ if $q(d^{\perp}-\gamma_{2})\equiv a_{1}\bmod q^{2}-1$, and $(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}$ if $\gamma_{0}=d^{\perp}-\gamma_{1}-\gamma_{2}>0$.
In the first case, $q(d^{\perp}-\gamma_{2})\equiv a_{1}\bmod q^{2}-1$ implies that $qd^{\perp}\equiv d\bmod q^{2}-1$, which happens if and only if $d\equiv 0\bmod q-1$. Taking into account that $d\leq 2(q-1)$, we have $d^{\perp}>q^{2}-1\geq\gamma_{2}$, and we have $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv(x_{1}^{d^{\perp}-\gamma_{2}}x_{2}^{\gamma_{2}})^{q}\bmod I(P^{2})$. Moreover, in this situation we can do this for all the monomials from $A_{2}^{d}$ in the support of $f$.
If $d\not\equiv 0\bmod q-1$, the only monomial in $(A^{d^{\perp}})^{q}$ with $x_{1}^{a_{1}}x_{2}^{a_{2}}$ in its expression in terms of the elements of the basis from Lemma 2.9 is $(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}$ with $\gamma_{0}=d^{\perp}-\gamma_{1}-\gamma_{2}$, if $d^{\perp}-\gamma_{1}-\gamma_{2}>0$. This is because, by Lemma 2.12, we have that
$(x_{0}^{\gamma_{0}}x_{1}^{\gamma_{1}}x_{2}^{\gamma_{2}})^{q}\equiv x_{1}^{a_{1}}x_{2}^{a_{2}}+(x_{0}-1)(x_{2}^{a_{2}}+x_{1}-1)\bmod I(P^{2}),$
@@
Arguing as in Lemma 3.5 and using Lemma 3.17, we obtain that there must be a sum of monomials in $((A^{d^{\perp}})^{q})^{(l)}$ equal to $\lambda_{0,d}\left((x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}\right)$. This implies that we have the evaluation of $(x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}$ in $\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},2)$. This is a contradiction because we have the evaluation of $(x_{0}^{d^{\perp}-\beta_{2}}x_{2}^{\beta_{2}})^{q}$ in $\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},2)$, and $(x_{0}-1)(x_{1}-1)+x_{0}x_{2}^{d}-(x_{0}^{d^{\perp}-\beta_{2}}x_{2}^{\beta_{2}})^{q}\equiv(x_{0}-1)(x_{1}-1)\bmod I(P^{2})$, whose evaluation has Hamming weight 1. ∎
@@
To state the next result, we use the following sets of polynomials. We recall that $U=\{x_{0}^{d-a_{1}-a_{2}}x_{1}^{a_{1}}x_{2}^{a_{2}}\mid x_{1}^{a_{1}}x_{2}^{a_{2}}\in U_{d-1,d}\}$, where we consider $U_{d-1,d}$ as in Definition 3.14. We define
$V:=\{x_{1}^{d-a_{2}}x_{2}^{a_{2}}\mid a_{2}\in T\},$
where $T$ is defined as in Lemma 3.20. Finally, for the case $12(q-1)<d\leq q^{2}-$ 1 we define
$W:=\{x_{1}^{d-a_{2}}x_{2}^{a_{2}}+x_{0}^{d-\overline{qd^{\perp}-a_{2}}-a_{2}}x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}\mid$ $q^{2}-1\geq d^{\perp}-\overline{qa_{2}}>\overline{q(d-a_{2})}$
[…]
We are interested in $W$ because of the following result.
Lemma 3.22.
We have $W\subset S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$, and $U\cup V\cup W$ is a linearly independent set in $S/I(P^{2})$.
Proof.
If $d^{\perp}-\overline{qa_{2}}>\overline{q(d-a_{2})}$ and $d-a_{2}>\overline{qd^{\perp}-a_{2}}$ (these conditions come from the definition of $W$), then
@@
This equivalence can be checked by considering the evaluation of both polynomials at $P^{2}$. Therefore, we have $W\subset S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. The condition $q^{2}-1\geq d^{\perp}-\overline{qa_{2}}$ ensures $U\cup V\cup W$ is linearly independent. Indeed, both of the monomials in the left hand side of (16) are monomials from the basis of Lemma 2.7 and are not in $U\cup V$. The monomial $x_{1}^{d-a_{2}}x_{2}^{a_{2}}$ from (16) is not in $V$ because $d^{\perp}\leq\overline{qa_{2}}-(q^{2}-1)$ (see the definition of $T$ in Lemma 3.20). For the monomial $x_{0}^{d-\overline{qd^{\perp}-a_{2}}-a_{2}}x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}$, it is not in $U$ because we have $x_{1}^{\overline{qd^{\perp}-a_{2}}}x_{2}^{a_{2}}\not\in U_{d-1,d}$. To see this, we use the definition of $U_{d-1,d}$ from Definition 3.14. First, $d-a_{2}>\overline{qd^{\perp}-a_{2}}$ means that this monomial satisfies the first condition in the definition of $U_{d-1,d}$. For the second one, we would have to check if
(17) $\overline{q(qd^{\perp}-a_{2})}+\overline{qa_{2}}=\overline{d^{\perp}-\overline{qa_{2}}}+\overline{qa_{2}}\leq d^{\perp}-1.$
If $\overline{d^{\perp}-\overline{qa_{2}}}<d^{\perp}-\overline{qa_{2}}$, we clearly have (17). However, if $\overline{d^{\perp}-\overline{qa_{2}}}=d^{\perp}-\overline{qa_{2}}$, we do not have (17), and this happens if and only if $q^{2}-1\geq d^{\perp}-\overline{qa_{2}}$, which is the condition we are using in the definition of $W$. ∎
Now we can state the main result of this section.
@@
Let $11\leq d\leq q^{2}-$ 1, and $U=\{x_{0}^{d-a_{1}-a_{2}}x_{1}^{a_{1}}x_{2}^{a_{2}}\mid x_{1}^{a_{1}}x_{2}^{a_{2}}\in U_{d-1,d}\}$, where we consider $U_{d-1,d}$ as in Definition 3.14. Let $V$ and $W$ be as in the discussion above. If $d\equiv 0\bmod q-1$, then $U\cup A_{2}^{d}\cup A_{3}^{d}$ is a basis for $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. For $d\not\equiv 0\bmod q-1$, if $d\leq 2(q-1)$, then $U\cup V=A_{1}^{d}\cup V$ is a basis for $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Lastly, if $12(q-1)<d\leq q^{2}-$ 1, then $U\cup V\cup W$ is a linearly independent set contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$.
@@
If $d\equiv 0\bmod q-1$ and $d\leq q^{2}-1$, reasoning as in the proofs of Lemmas 3.17, 3.20 and 3.21, we have $U\cup A_{2}^{d}\cup A_{3}^{d}$ contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Let $f\in S_{d}$ such that $f\in S_{d^{\perp}}^{q}/I(P^{2})$ and whose support is contained in $A_{1}^{d}\setminus U$. Then $f$ has an expression in terms of the monomials from $(A^{d^{\perp}})^{q}$ in $S/I(P^{2})$, and this expression only involves the monomials from $(A_{1}^{d^{\perp}})^{q}$. This is because $x_{2}^{qd^{\perp}}$ cannot be in the expression because it is nonzero at $[0:0:1]$ while $f(0,0,1)=0$, and if there were monomials from $(A_{2}^{d^{\perp}})^{q}$, in [$[\{0\}\times\{1\}\times{\mathbb{F}}_{q^{2}}]$ that expression would have the same evaluation as some polynomial in $x_{2}$ with degree less than or equal to $q^{2}-1$, which cannot have $q^{2}$ zeroes ($f$ is equal to 00 in those points). We finish this case by using the affine case from Proposition 3.15.
@@
Finally, in the case $d\not\equiv 0\bmod q-1$ and $12(q-1)<d<q^{2}-$ 1, we have from Lemma 3.17 that $U$ is contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. The fact that $V$ is contained in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ follows from (12). For $W$, by Lemma 3.22, we have that $W\subset S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ and $U\cup V\cup W$ is linearly independent in $S/I(P^{2})$. ∎
From Theorem 3.23 we can obtain the exact dimension of $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ if $d\leq 2(q-1)$ or $d\equiv 0\bmod q-1$. For $12(q-1)<d\leq q^{2}-$ 1, $d\not\equiv 0\bmod q-1$, we only have a lower bound for the dimension of $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$. Note that a lower bound for the dimension of the hull gives an upper bound for the parameter $c$ of the corresponding EAQECC, which is still interesting because it tells us how many maximally entangled pairs are required at most for using that EAQECC. Equivalently, if we use as many maximally entangled pairs as the bound specifies, then we can employ this EAQECC. Nevertheless, in all cases we have checked, this is indeed the true value of the dimension of the hull, which implies that $U\cup V\cup W$ is also a basis for the hull in those cases. In particular, this means that the Hermitian hull is not generated by monomials in general because of $W$ (we saw in the proof of Lemma 3.22 that no monomial of the polynomials from $W$ is contained in $U\cup V$). We see this in the next example.
@@ 4.2. Hermitian EAQECCs
Let $10\leq d<q^{2}-$ 1. Then we can construct an EAQECC with parameters $[[n,\kappa,\delta;c]]_{q}$, where $n=q^{4}$, $\kappa=n-2(\dim\operatorname{RM}_{d}(q^{2},2))+c$, $\delta\geq\operatorname{wt}(\operatorname{RM}_{d}^{\perp}(q^{2},2))$, and the value of $c$ is